The Ship Corp. has paid annual dividends of RM0.48, RM0.60, and RM0.62 a share over the
past three years, respectively. The company now predicts that it will maintain a constant
dividend since its business has leveled off and sales are expected to remain relatively
constant. Given the lack of future growth, at what price will you only buy this stock if you
can earn at least a 14 percent rate of return?
1
Expert's answer
2016-05-07T12:17:03-0400
The Ship Corp. has paid annual dividends of RM0.48, RM0.60, and RM0.62 a share over the past three years, respectively. The company now predicts that it will maintain a constant dividend since its business has leveled off and sales are expected to remain relatively constant, rate of return is 14%.
A constant growth stock is a stock whose dividends are expected to grow at a constant rate in the forseeable future. This condition fits many established firms, which tend to grow over the long run at the same rate as the economy, fairly well. The value of a constant growth stock can be determined using the following equation:
P0 = D0*(1 + g)/(r - g) = D1/(r - g), where P0 = the stock price at time 0, D0 = the current dividend, D1 = the next dividend (i.e., at time 1), g = the growth rate in dividends, and r = the required return on the stock, and g < r.
g = D2/D1 - 1 = 0.62/0.6 - 1 = 0.033.
So, P0 = 0.62*(1 + 0.033)/(0.14 - 0.033) = $5.99 per share.